### 21 Dec 2023 v5

@@ 2.1. Proof systems
Definition 2.1 ([CCT87]).
The Cutting Planes (CP) proof system is equipped with boolean axioms and two inference rules:
[…]
@@ 3.2. Sperner’s Theorem
We consider the following extension of Sperner’s theorem.
Theorem 4 ([MR08, CCT09]).
Fix any $t\geq 2,t\in\mathbb{N}$. For all $f\in\mathbb{N}$, with the pointwise ordering of $[t]^{f}$, any antichain has size at most $t^{f}\sqrt{\frac{6}{\pi(t^{2}-1)f}}(1+o(1)).$
@@
Define $I_{\mathbf{a}}=\{i\in[n]:a_{i}\not=0\}$. Let $\preceq$ be the partial order over $W^{I_{\mathbf{a}}}$ where $\mathbf{x}\preceq\mathbf{y}$ if $x_{i}\leq y_{i}$ for all $i$ with $a_{i}>0$ and $x_{i}\geq y_{i}$ for the remaining $i$ with $a_{i}<0$. Clearly the set of solutions to $\mathbf{a}\mathbf{s}=b$ forms an antichain under $\preceq$. Noting that $\preceq$ is isomorphic to the typical pointwise ordering on $W^{I_{\mathbf{a}}}$, we appeal to Theorem 4 to upper bound the number of solutions in $W^{I_{\mathbf{a}}}$ by $\frac{k^{w(\mathbf{a})}}{\sqrt{w(\mathbf{a})}}$, each of which corresponds to at most $k^{n-w(\mathbf{a})}$ vectors in $W^{n}$.
@@ 3.3. Large admissibility
Since all the queries in ${\mathcal{T}}$ have width at least $w$, according to Lemma 5, each query in ${\mathcal{T}}$ rules out at most $\ell\cdot\frac{k^{\nu(n)}}{\sqrt{w}}$ admissible points. By Fact 1 no point survives at the leaves, in particular the admissible points. Then it must be that
@@ 3.4. Main theorem
holds. At the stage $i$ we let $\Sigma_{i}=\{(\mathbf{a},b)\in{\mathcal{T}}_{i}:w(\mathbf{a})\leq t^{2}\}$ and $s_{i}=|\Sigma_{i}|$. If $s_{i}\geq t$ the claim is trivially proven. If $s_{i}=0$, then all queries in ${\mathcal{T}}_{i}$ have width at least $t^{2}$ and by Lemma 7 (which can be applied since $n-ict^{3}>n_{{\mathscr{F}}}$) the claim is proven (for $n$ large enough). So assume that $0<s_{i}<t$. Each of the queries in $\Sigma_{i}$ involves at most $t^{2}$ nonzero coefficients, hence in total they mention at most $s_{i}t^{2}\leq t^{3}$ variables. Extend this set of variables to some $V^{\prime}$ in accordance with Definition 3.9 (which can be done since, by the invariant, $ict^{3}<n/c$). Set all these variables according to self-reducibility of ${\mathcal{F}}$ in a restriction $\rho_{i}$ and define ${\mathcal{T}}_{i+1}={\mathcal{T}}_{i}\!\!\restriction\!\!_{\rho_{i}}$. Note that by Definition 3.9 and by that of restriction, ${\mathcal{T}}_{i+1}$ is a $\operatorname{\mathsf{SP}}$ refutation of ${\mathcal{F}}_{n-ict^{3}}$ and we can go on with the next stage. (Also note that we do not hit an empty refutation this way, due to the assumption that ${\mathcal{F}}_{n}$ defines a feasible LP.) Assume that the invariant does not hold. If this is because $n-ict^{3}<n_{{\mathscr{F}}}$ then, as each iteration destroys at least one node,
@@
Using Lemmas 6 and 8 and the previous Theorem we get:
@@ 3.5. Lower bounds for the Pigeonhole principle
Note that the all $1/2$ point is feasible for $\operatorname{\mathsf{PHP}}^{n+d}_{n}$. Then with Lemma 11 and Lemma 12 in hand we meet all the prerequisites for Theorem 9.
@@ 3.6. Lower bounds for Tseitin contradictions over the complete graph
We have that the all $1/2$ point is feasible for $\operatorname{\mathsf{Ts}}(K_{n},\omega)$. Then we can simply apply Theorem 9.
@@ 3.7. Lower bound for the Least Ordering Principle
We will assume that a $\operatorname{\mathsf{SP}}$ refutation ${\mathcal{T}}$ of $\operatorname{\mathsf{LOP}}_{n}$ only involves variables $P_{i,j}$ where $i<j$ this is without loss of generality as we can safely set $P_{j,i}$ to $1-P_{i,j}$ whenever $i>j$, and will often write $P_{\{i,j\}}$ for such a variable. We consider the underlying graph of the support of a query, i.e. an undirected graph with edges $\{i,j\}$ for every variable $P_{\{i,j\}}$ that appears with non-zero coefficient in the query. For some function $f(n)$, we say the query is $f(n)$-wide if the smallest edge cover of its graph has at least $f(n)$ nodes . A query that is not $f(n)$-wide is $f(n)$-narrow. The next lemma works much the same as Theorem 9.
@@
We iteratively build up an initially empty restriction $\rho$. At every stage $\rho$ imposes a total order on some subset $X\subseteq[n]$ and places the elements in $X$ above the elements not in $X$. So $\rho$ sets every edge not contained entirely in $[n]\setminus X$ to something binary, and $\operatorname{\mathsf{LOP}}_{n}\!\!\restriction_{\rho}=\operatorname{\mathsf{LOP}}_{n-|X|}$ (up to a renaming of variables). While there exists a $n^{3/4}$-narrow query $q\in{\mathcal{T}}\!\!\restriction_{\rho}$ we simply take its smallest edge cover, which has size at most $n^{3/4}$ by definition, and add its nodes in any fashion to the total order in $\rho$. Now all of the variables mentioned by $q\in{\mathcal{T}}\!\!\restriction_{\rho}$ are fully evaluated and $q$ is redundant. We repeat this at most $n^{\frac{1-\epsilon}{4}}$ times (as $|{\mathcal{T}}|\leq n^{\frac{1-\epsilon}{4}}$ and each iteration renders at least one query in ${\mathcal{T}}$ redundant). At each stage we grow the domain of the restriction by at most $n^{3/4}$, so the domain of $\rho$ is always bounded by $n^{1-\epsilon/4}$. We also cannot exhaust the tree ${\mathcal{T}}$ in this way, as otherwise ${\mathcal{T}}$ mentioned at most $n^{1-\epsilon/4}<n-3$ elements and by Lemma 19 there is an admissible point not falling in any slab of ${\mathcal{T}}$, violating Fact 1.
@@
Suppose otherwise - then, by Lemma 20, we can find some ${\mathcal{T}}^{\prime}$ refuting $\operatorname{\mathsf{LOP}}_{cn}$, with $|{\mathcal{T}}^{\prime}|\leq|{\mathcal{T}}|$, every query $n^{3/4}$-wide, and $c$ independent of $n$. We greedily create a set of pairs $D$ by processing the queries in ${\mathcal{T}}^{\prime}$ one by one and choosing in each a matching of size $n^{1/2}$ disjoint from the elements appearing in $D$ this always succeeds, as at every stage $|D|\in O(n^{\frac{1-\epsilon}{4}}\cdot n^{1/2})$ and involves at most $O(2n^{\frac{3-\epsilon}{4}})<n^{3/4}-n^{1/2}$ elements.
So by Lemma 21, after setting every edge not in $D$ to $1/2$, we have some set of linear polynomials =${\mathcal{R}}=\{a(x)=\mathbf{a}x-b-1/2:(\overline{a},b)\in{\mathcal{T}}^{\prime}\}$ covering the hypercube $\{0,1\}^{D}$, where every polynomial $p\in{\mathcal{R}}$ mentions at least $n^{1/2}$ edges. By Lemma 5 each such polynomial in ${\mathcal{R}}$ rules out at most $\nicefrac{{2^{|D|}}}{{n^{1/4}}}$ points, and so we must have $|{\mathcal{T}}|\geq|T^{\prime}|\geq|{\mathcal{R}}|\geq n^{1/4}$.
@@ 4. The covering method
Definition 4.2 ([LR05]).
A set $L$ of linear polynomials with real coefficients is said to be an essential cover of the cube $\{0,1\}^{n}$ if
@@ 4.1. The covering method and Tseitin
Let $H_{n}$ denote the $n\times n$ grid graph. Fix some $\omega$ with odd charge and a $\operatorname{\mathsf{SP}}$ refutation ${\mathcal{T}}$ of $\operatorname{\mathsf{Ts}}(H_{n},\omega)$. Fact 1 tells us that for every point $x$ admissible for $\operatorname{\mathsf{Ts}}(H_{n},\omega)$, there exists a query $(\mathbf{a},b)\in{\mathcal{T}}$ such that $b<\mathbf{a}x<b+1$. In this section we will only consider admissible points with entries in $\{0,1/2,1\}$, turning the slab of a query $(\mathbf{a},b)$ into the solution set of the single linear equation $\mathbf{a}\cdot x=b+1/2$. So we consider ${\mathcal{T}}$ as a set of such equations.
@@
(where $b_{s}$ is from Lemma 25).
@@
Proof 4.8.
When $\alpha(v_{t})$ is all 00, $h_{s}=b_{s}$ is admissible (by Lemma 25). Toggling some $v_{t}$ only has the effect of flipping every edge in a cycle, which preserves admissibility.
@@
As $|S|=(n/3)^{2}$ this would imply that $\sum_{i=1}^{q}|M_{i}|\geq(n/3)^{2}-q$. If $q\geq(n/3)^{2}/2$, then (as the $L_{i}$ are nonempty and pairwise disjoint) we have $|{\mathcal{T}}|\geq(n/3)^{2}/2\in\Omega(n^{1.04})$. Otherwise $\sum_{i=1}^{q}|M_{i}|\geq(n/3)^{2}/2$, and as (by Theorem 23) each $|L_{i}|\geq|M_{i}|^{0.52}$,
@@
We create the pairs by stages. Let $S_{1}=S$ and start by picking any $s_{1}\in S_{1}$. By Lemma 28 ${\mathcal{T}}_{s_{1}}$ covers $\{0,1\}^{V_{S_{1}}\setminus\{s_{1}\}}$ and has as an essentialisation $E$, which will be an essential cover of $\{0,1\}^{V^{\prime}}$ for some $V^{\prime}\subseteq V_{S_{1}}\setminus\{s_{1}\}$. We create the pair $(L_{1},M_{1})=(\{p:p_{s_{1}}\in E\},V^{\prime})$ and update $S_{2}=S_{1}\setminus\left(V^{\prime}\cup\{s_{1}\}\right)$. (Note that $V^{\prime}$ could possibly be empty - for example, if the polynomial $x_{e}=1/2$ appears in ${\mathcal{T}}$, where $e\in s_{1}$. In this case however we still have $|L_{1}|\geq|M_{1}|^{0.52}$. If $V^{\prime}$ is not empty we have the same bound due to Theorem 23.) If $S_{2}$ is nonempty we repeat with any $s_{2}\in S_{2}$, and so on. We now show that as promised the left hand sides of these pairs partition a subset of ${\mathcal{T}}$, which will give us the first inequality in Equation 2. Every polynomial $p$ with $p_{s_{i}}\in L_{i}$ has every $v_{t}$ mentioned by $p_{s_{i}}$ removed from $S_{j}$ for all $j\geq i$, so the only way $p$ could reappear in some later $L_{j}$ is if $p_{s_{j}}\in{\mathcal{T}}_{s_{j}}$, where $v_{s_{j}}$ does not appear in $p_{s_{i}}$. Let $\mu_{e},e\in s_{j}$ be the coefficients of $p$ in front of the four edges of $s_{j}$. The coefficient in front of $v_{s_{j}}$ in $p_{s_{i}}$ is just $\sum_{e\in s_{j}}\mu_{e}$. As $v_{s_{j}}$ failed to appear this sum is 00 and $p$ does not have the odd coefficient sum it would need to appear in ${\mathcal{T}}_{s_{j}}$.

### 10 Jan 2024 v6

@@ 2.1. Proof systems
{defiC}
[[CCT87]] The Cutting Planes (CP) proof system is equipped with boolean axioms and two inference rules:
Boolean AxiomsLinear $\begin{array}[]{c|c|c}\mbox{Boolean Axioms}&\mbox{Linear Combination}&\mbox{Rounding}\\ \frac{\quad\quad}{x\geq 0}\quad\frac{\quad\quad}{-x\geq-1}&\frac{\mathbf{a}\mathbf{x}\geq c\quad\quad\mathbf{b}\mathbf{x}\geq d}{\alpha\mathbf{a}\mathbf{x}+\beta\mathbf{b}\mathbf{x}\geq\alpha c+\beta d}&\frac{\alpha\mathbf{a}\mathbf{x}\geq b}{\mathbf{a}\mathbf{x}\geq\lceil b/\alpha\rceil}\end{array}$
@@ 3.2. Sperner’s Theorem
We consider the following extension of Sperner’s theorem. {thmC}[[MR08, CCT09]] Fix any $t\geq 2,t\in\mathbb{N}$. For all $f\in\mathbb{N}$, with the pointwise ordering of $[t]^{f}$, any antichain has size at most $t^{f}\sqrt{\frac{6}{\pi(t^{2}-1)f}}(1+o(1)).$
@@
Define $I_{\mathbf{a}}=\{i\in[n]:a_{i}\not=0\}$. Let $\preceq$ be the partial order over $W^{I_{\mathbf{a}}}$ where $\mathbf{x}\preceq\mathbf{y}$ if $x_{i}\leq y_{i}$ for all $i$ with $a_{i}>0$ and $x_{i}\geq y_{i}$ for the remaining $i$ with $a_{i}<0$. Clearly the set of solutions to $\mathbf{a}\mathbf{s}=b$ forms an antichain under $\preceq$. Noting that $\preceq$ is isomorphic to the typical pointwise ordering on $W^{I_{\mathbf{a}}}$, we appeal to Definition 3.3 to upper bound the number of solutions in $W^{I_{\mathbf{a}}}$ by $\frac{k^{w(\mathbf{a})}}{\sqrt{w(\mathbf{a})}}$, each of which corresponds to at most $k^{n-w(\mathbf{a})}$ vectors in $W^{n}$.
@@ 3.3. Large admissibility
Since all the queries in ${\mathcal{T}}$ have width at least $w$, according to Lemma 3.2, each query in ${\mathcal{T}}$ rules out at most $\ell\cdot\frac{k^{\nu(n)}}{\sqrt{w}}$ admissible points. By Fact 2.1 no point survives at the leaves, in particular the admissible points. Then it must be that
@@ 3.4. Main theorem
holds. At the stage $i$ we let $\Sigma_{i}=\{(\mathbf{a},b)\in{\mathcal{T}}_{i}:w(\mathbf{a})\leq t^{2}\}$ and $s_{i}=|\Sigma_{i}|$. If $s_{i}\geq t$ the claim is trivially proven. If $s_{i}=0$, then all queries in ${\mathcal{T}}_{i}$ have width at least $t^{2}$ and by Lemma 3.3 (which can be applied since $n-ict^{3}>n_{{\mathscr{F}}}$) the claim is proven (for $n$ large enough). So assume that $0<s_{i}<t$. Each of the queries in $\Sigma_{i}$ involves at most $t^{2}$ nonzero coefficients, hence in total they mention at most $s_{i}t^{2}\leq t^{3}$ variables. Extend this set of variables to some $V^{\prime}$ in accordance with Definition 3.9 (which can be done since, by the invariant, $ict^{3}<n/c$). Set all these variables according to self-reducibility of ${\mathcal{F}}$ in a restriction $\rho_{i}$ and define ${\mathcal{T}}_{i+1}={\mathcal{T}}_{i}\!\!\restriction\!\!_{\rho_{i}}$. Note that by Definition 3.9 and by that of restriction, ${\mathcal{T}}_{i+1}$ is a $\operatorname{\mathsf{SP}}$ refutation of ${\mathcal{F}}_{n-ict^{3}}$ and we can go on with the next stage. (Also note that we do not hit an empty refutation this way, due to the assumption that ${\mathcal{F}}_{n}$ defines a feasible LP.) Assume that the invariant does not hold. If this is because $n-ict^{3}<n_{{\mathscr{F}}}$ then, as each iteration destroys at least one node,
@@
Using Lemmas 3.3 and 3.4 and the previous Theorem we get:
@@ 3.5. Lower bounds for the Pigeonhole principle
Note that the all $1/2$ point is feasible for $\operatorname{\mathsf{PHP}}^{n+d}_{n}$. Then with Section 3.5 and Section 3.5 in hand we meet all the prerequisites for Section 3.4.
@@ 3.6. Lower bounds for Tseitin contradictions over the complete graph
We have that the all $1/2$ point is feasible for $\operatorname{\mathsf{Ts}}(K_{n},\omega)$. Then we can simply apply Section 3.4.
@@ 3.7. Lower bound for the Least Ordering Principle
We will assume that a $\operatorname{\mathsf{SP}}$ refutation ${\mathcal{T}}$ of $\operatorname{\mathsf{LOP}}_{n}$ only involves variables $P_{i,j}$ where $i<j$ this is without loss of generality as we can safely set $P_{j,i}$ to $1-P_{i,j}$ whenever $i>j$, and will often write $P_{\{i,j\}}$ for such a variable. We consider the underlying graph of the support of a query, i.e. an undirected graph with edges $\{i,j\}$ for every variable $P_{\{i,j\}}$ that appears with non-zero coefficient in the query. For some function $f(n)$, we say the query is $f(n)$-wide if the smallest edge cover of its graph has at least $f(n)$ nodes . A query that is not $f(n)$-wide is $f(n)$-narrow. The next lemma works much the same as Section 3.4.
@@
We iteratively build up an initially empty restriction $\rho$. At every stage $\rho$ imposes a total order on some subset $X\subseteq[n]$ and places the elements in $X$ above the elements not in $X$. So $\rho$ sets every edge not contained entirely in $[n]\setminus X$ to something binary, and $\operatorname{\mathsf{LOP}}_{n}\!\!\restriction_{\rho}=\operatorname{\mathsf{LOP}}_{n-|X|}$ (up to a renaming of variables). While there exists a $n^{3/4}$-narrow query $q\in{\mathcal{T}}\!\!\restriction_{\rho}$ we simply take its smallest edge cover, which has size at most $n^{3/4}$ by definition, and add its nodes in any fashion to the total order in $\rho$. Now all of the variables mentioned by $q\in{\mathcal{T}}\!\!\restriction_{\rho}$ are fully evaluated and $q$ is redundant. We repeat this at most $n^{\frac{1-\epsilon}{4}}$ times (as $|{\mathcal{T}}|\leq n^{\frac{1-\epsilon}{4}}$ and each iteration renders at least one query in ${\mathcal{T}}$ redundant). At each stage we grow the domain of the restriction by at most $n^{3/4}$, so the domain of $\rho$ is always bounded by $n^{1-\epsilon/4}$. We also cannot exhaust the tree ${\mathcal{T}}$ in this way, as otherwise ${\mathcal{T}}$ mentioned at most $n^{1-\epsilon/4}<n-3$ elements and by Section 3.7 there is an admissible point not falling in any slab of ${\mathcal{T}}$, violating Section 2.1.
@@
Suppose otherwise - then, by Section 3.7, we can find some ${\mathcal{T}}^{\prime}$ refuting $\operatorname{\mathsf{LOP}}_{cn}$, with $|{\mathcal{T}}^{\prime}|\leq|{\mathcal{T}}|$, every query $n^{3/4}$-wide, and $c$ independent of $n$. We greedily create a set of pairs $D$ by processing the queries in ${\mathcal{T}}^{\prime}$ one by one and choosing in each a matching of size $n^{1/2}$ disjoint from the elements appearing in $D$ this always succeeds, as at every stage $|D|\in O(n^{\frac{1-\epsilon}{4}}\cdot n^{1/2})$ and involves at most $O(2n^{\frac{3-\epsilon}{4}})<n^{3/4}-n^{1/2}$ elements.
So by Section 3.7, after setting every edge not in $D$ to $1/2$, we have some set of linear polynomials =${\mathcal{R}}=\{a(x)=\mathbf{a}x-b-1/2:(\overline{a},b)\in{\mathcal{T}}^{\prime}\}$ covering the hypercube $\{0,1\}^{D}$, where every polynomial $p\in{\mathcal{R}}$ mentions at least $n^{1/2}$ edges. By Section 3.2 each such polynomial in ${\mathcal{R}}$ rules out at most $\nicefrac{{2^{|D|}}}{{n^{1/4}}}$ points, and so we must have $|{\mathcal{T}}|\geq|T^{\prime}|\geq|{\mathcal{R}}|\geq n^{1/4}$.
@@ 4. The covering method
{defiC}
[[LR05]] A set $L$ of linear polynomials with real coefficients is said to be an essential cover of the cube $\{0,1\}^{n}$ if
@@ 4.1. The covering method and Tseitin
Let $H_{n}$ denote the $n\times n$ grid graph. Fix some $\omega$ with odd charge and a $\operatorname{\mathsf{SP}}$ refutation ${\mathcal{T}}$ of $\operatorname{\mathsf{Ts}}(H_{n},\omega)$. Section 2.1 tells us that for every point $x$ admissible for $\operatorname{\mathsf{Ts}}(H_{n},\omega)$, there exists a query $(\mathbf{a},b)\in{\mathcal{T}}$ such that $b<\mathbf{a}x<b+1$. In this section we will only consider admissible points with entries in $\{0,1/2,1\}$, turning the slab of a query $(\mathbf{a},b)$ into the solution set of the single linear equation $\mathbf{a}\cdot x=b+1/2$. So we consider ${\mathcal{T}}$ as a set of such equations.
@@
(where $b_{s}$ is from Section 4.1).
@@
Proof 4.7.
When $\alpha(v_{t})$ is all 00, $h_{s}=b_{s}$ is admissible (by Section 4.1). Toggling some $v_{t}$ only has the effect of flipping every edge in a cycle, which preserves admissibility.
@@
As $|S|=(n/3)^{2}$ this would imply that $\sum_{i=1}^{q}|M_{i}|\geq(n/3)^{2}-q$. If $q\geq(n/3)^{2}/2$, then (as the $L_{i}$ are nonempty and pairwise disjoint) we have $|{\mathcal{T}}|\geq(n/3)^{2}/2\in\Omega(n^{1.04})$. Otherwise $\sum_{i=1}^{q}|M_{i}|\geq(n/3)^{2}/2$, and as (by Section 4) each $|L_{i}|\geq|M_{i}|^{0.52}$,
@@
We create the pairs by stages. Let $S_{1}=S$ and start by picking any $s_{1}\in S_{1}$. By Section 4.1 ${\mathcal{T}}_{s_{1}}$ covers $\{0,1\}^{V_{S_{1}}\setminus\{s_{1}\}}$ and has as an essentialisation $E$, which will be an essential cover of $\{0,1\}^{V^{\prime}}$ for some $V^{\prime}\subseteq V_{S_{1}}\setminus\{s_{1}\}$. We create the pair $(L_{1},M_{1})=(\{p:p_{s_{1}}\in E\},V^{\prime})$ and update $S_{2}=S_{1}\setminus\left(V^{\prime}\cup\{s_{1}\}\right)$. (Note that $V^{\prime}$ could possibly be empty - for example, if the polynomial $x_{e}=1/2$ appears in ${\mathcal{T}}$, where $e\in s_{1}$. In this case however we still have $|L_{1}|\geq|M_{1}|^{0.52}$. If $V^{\prime}$ is not empty we have the same bound due to Section 4.) If $S_{2}$ is nonempty we repeat with any $s_{2}\in S_{2}$, and so on. We now show that as promised the left hand sides of these pairs partition a subset of ${\mathcal{T}}$, which will give us the first inequality in Equation 2. Every polynomial $p$ with $p_{s_{i}}\in L_{i}$ has every $v_{t}$ mentioned by $p_{s_{i}}$ removed from $S_{j}$ for all $j\geq i$, so the only way $p$ could reappear in some later $L_{j}$ is if $p_{s_{j}}\in{\mathcal{T}}_{s_{j}}$, where $v_{s_{j}}$ does not appear in $p_{s_{i}}$. Let $\mu_{e},e\in s_{j}$ be the coefficients of $p$ in front of the four edges of $s_{j}$. The coefficient in front of $v_{s_{j}}$ in $p_{s_{i}}$ is just $\sum_{e\in s_{j}}\mu_{e}$. As $v_{s_{j}}$ failed to appear this sum is 00 and $p$ does not have the odd coefficient sum it would need to appear in ${\mathcal{T}}_{s_{j}}$.